### v1

@@ Abstract
We study linear regression models with clustered data, high-dimensional controls, and a complicated structure of exclusion restrictions. We propose a correctly centered internal IV estimator that accommodates a variety of exclusion restrictions and permits within-cluster dependence. The estimator has a simple leave-out interpretation and remains computationally tractable. We derive a central limit theorem for its quadratic form and propose a robust variance estimator. We also develop inference methods that remain valid under weak identification. Our framework extends classical dynamic panel methods to more general clustered settings. An empirical application of a large-scale fiscal intervention in rural Kenya with spatial interference illustrates the approach.
Keywords: weak exogeneity, many controls, feedback bias, OLS inconsistency, internal instrument, leave-out approach, valid inference
@@ 1 Introduction
Clustered data—such as panel, network, spatial, or group-structured datasets, where individuals or firms are observed repeatedly or nested within groups—have become increasingly common in empirical research. These settings present several methodological challenges. One is the need to account for complex heterogeneity, often addressed through the inclusion of fixed effects and flexible time or group trends, which typically entails incorporating a large number of control variables. A more fundamental difficulty arises from within-cluster dependence: observations within the same cluster may exhibit correlation structures, including spatial and network interference, spillover effects, and time-series dependence. Such dependencies complicate statistical inference and limit the effective aggregation of information across observations.
@@
The concept of exogeneity becomes more nuanced in the presence of clustered data. As we demonstrate, assuming only a per-observation exclusion restriction—i.e., that the error term is uncorrelated with the regressor within the same observation—may fail to provide any identifying variation for consistent estimation of structural parameters. Conversely, imposing strict exogeneity, where each error term is uncorrelated with all regressors in the same cluster, is often too strong and implausible in many empirical contexts.
@@
To bridge the gap between strict exogeneity and unrestricted dependence, we adopt an intermediate assumption: the error term is uncorrelated with a subset of regressors from the same cluster. The form of this exclusion restriction depends on the application and institutional context. In panel data, it is common to assume errors are mean-zero conditional on current and past—but not future—regressors, allowing for feedback from shocks to future policies. In spatial settings, localized spillovers justify similar structures—for example, spatial leakage in deforestation (Jayachandran et al., 2017), spillovers in policing (Blattman et al., 2021), or a fiscal intervention experiment in rural Kenya (Egger et al., 2022). In networks, treatment effects may propagate through social ties (Paluck et al., 2016). These partial exogeneity assumptions better reflect the complex dependence seen in many empirical settings.
@@
When we assume only a subset of exclusion restrictions, it becomes inappropriate to treat regressors as fixed or condition on them. Their stochastic variation may correlate with the errors, invalidating standard estimation approaches. As a result, even estimators like OLS must be interpreted as ratios of stochastic quadratic forms, introducing several challenges.
@@
The most well-known is Nickell bias (Nickell, 1981), which occurs when the expected value of the OLS error numerator is nonzero, leading to asymptotic bias. While prominent in dynamic panels with fixed effects and lagged outcomes, this bias arises more broadly under clustered dependence. A second challenge is inference: standard variance estimators often fail to account for the cross-cluster dependence of quadratic forms. Finally, the denominator may remain random asymptotically, unlike in classical settings, giving rise to weak identification and complicating inference further.
@@
This paper studies the estimation of structural parameters in linear regression models with clustered data, high-dimensional controls, and a researcher-specified set of exclusion restrictions. A key feature of the framework is the grouping of observations into disjoint clusters, permitting within-cluster dependence and independence across clusters. The setting accommodates unbalanced panels with multiple high-dimensional fixed effects, as well as spatial, network, and group-structured data. Our approach extends classical dynamic panel estimators to a broader class of models, addressing the central challenges of bias, inference, and identification.
@@
Our first contribution is to characterize a class of correctly centered internal instrument estimators that eliminate the asymptotic bias of OLS. The approach accommodates high-dimensional exogenous controls and adapts to the exclusion structure specified in the application. The resulting estimator is an internal IV estimator, defined as the one closest to OLS in a well-defined norm. Under some conditions, the estimator is asymptotically efficient and easy to implement.
@@
The procedure has a simple interpretation: for each observation, controls are partialled out using only those observations whose errors are uncorrelated with its regressor, yielding an observation-specific leave-out projection. A just-identified IV regression is then performed on the transformed equation, using the original regressor as an instrument.
@@
We further characterize the efficiency loss from varying the strength of exclusion restrictions. In particular, assuming exogeneity only within the same observation eliminates all identifying variation when fixed effects are present. This observation underscores the importance of carefully specifying the exogeneity structure in applied work.
@@
Our second contribution addresses inference. We show that, except in two special cases—(i) strict exogeneity and (ii) block-diagonal residualization (e.g., models with only cluster fixed effects)—the estimator’s numerator is a nontrivial quadratic form in the errors. In such settings, standard cluster-robust variance estimators may fail, and valid inference requires central limit theorems for clustered quadratic forms. These issues are especially acute in models with many high-dimensional controls, such as two-way fixed effects (Verdier, 2020). We propose a new central limit theorem and a new variance estimator that account for this structure.
@@
Our third contribution addresses weak identification, which arises when the internal instrument—though valid—captures limited identifying variation due to many controls or weak exclusion restrictions. As in the classical dynamic panel literature (Blundell and Bond, 1998; Bun and Windmeijer, 2010), weak identification can compromise standard inference. We develop inference procedures that remain valid even when the estimator’s denominator exhibits substantial sampling variability.
@@
We propose a comprehensive empirical strategy that combines a correctly centered estimator with identification-robust inference and confidence sets. We apply it to a prominent randomized evaluation of a large-scale fiscal intervention in rural Kenya (Egger et al., 2022). A key challenge in this setting is spatial interference: treatment in one village affects outcomes in neighboring villages, complicating the choice of plausible exclusion restrictions. We show how relaxing these assumptions leads to less precise estimates and wider confidence sets.
@@
This paper contributes to several strands of the econometrics literature. It relates to the extensive work on linear dynamic panel data models—a mature field with too many important contributions to list exhaustively. For recent overviews on correcting Nickell bias, we refer to the surveys by Okui (2021) and Bun and Sarafidis (2015). Our estimation approach is more closely aligned with the internal instrument framework developed in Anderson and Hsiao (1981), Arellano and Bond (1991), Ahn and Schmidt (1995), Alvarez and Arellano (2003), and Blundell and Bond (1998), though we extend the framework to accommodate more general models. In addition, we directly address weak identification concerns noted by the dynamic panel literature, particularly in Bun and Windmeijer (2010). We also contribute to the emerging literature on quadratic central limit theorems for clustered data structures. These results have been recently employed in many-instrument panel settings, such as Ligtenberg (2023), and are crucial for understanding the distribution of estimators that take the form of quadratic error functionals.
@@
The remainder of the paper is organized as follows. Section 2 describes the data structure, a plausible set of exclusion restrictions, and two modeling perspectives—outcome-based and design-based. Section 3 characterizes the class of correctly centered internal instrument estimators, introduces our proposed estimator as the solution to an efficiency optimization problem, and interprets it as a leave-out internal IV procedure. Section 4 addresses uncertainty quantification, and Section 5 establishes a central limit theorem for quadratic forms with clustered data. In Section 6, we develop inference procedures robust to weak identification and propose a corresponding variance estimator. Section 7 presents an empirical application that illustrates the full strategy.
@@
We use $0<c<1$ and $C>0$ to denote generic finite constants that may vary across equations but do not depend on the sample size. For a matrix $A$, we let $A^{+}$ denote its Moore–Penrose generalized inverse, $\operatorname{tr}(A)$ its trace, $\|A\|_{F}^{2}=\operatorname{tr}(A^{\prime}A)$ its squared Frobenius norm, and $\|A\|$ its operator (spectral) norm.
@@ 2.1 Data structure
where $\ell$ indexes the $n$ observations in the dataset. The parameter of interest, $\beta$, is a scalar, though most of our results extend to vectors of fixed dimension. The vector of strictly exogenous controls $w_{\ell}$ has dimension $K$, which may be large but is assumed to satisfy $K<n$. We denote by $W=[w_{1},\dots,w_{n}]^{\prime}$ the matrix of control variables. For simplicity, we treat $W$ as non-random, although our results can alternatively be interpreted as conditional on $W$. Throughout, we assume that $W$ has full rank and define the projection matrix $M=I_{n}-W(W^{\prime}W)^{-1}W^{\prime},$ which projects out the variation associated with the controls $W$.
@@
Consider a regression with individual and time fixed effects: $y_{it}=\alpha_{i}+\mu_{t}+\beta x_{it}+e_{it},$ where $y_{it}$ and $x_{it}$ denote the outcome and treatment for student $i$ at time $t$. Let $\ell=\ell(i,t)$ index the observations, and define clusters as $S_{i}=\{\ell(i,t):\text{for all observed }t\}$. Students may be observed in different and potentially non-overlapping periods. Clusters need not coincide with individual indices $i$; for example, one may define clusters at the family level if students are siblings. $\Box$
@@
Suppose we estimate the causal effect of a policy $x_{\ell}$, implemented in municipality $\ell$, on outcome $y_{\ell}$. Assume that municipalities are nested within metropolitan areas $i$, which define the clusters. A rich set of controls is included—such as cluster fixed effects and federally influenced covariates—so that $(x_{\ell},e_{\ell})$ are assumed independent across clusters, but may be arbitrarily dependent within clusters. $\Box$
@@
Consider treatment and outcome data for individuals connected via a social or informational network. The data may be drawn either from distinct clusters within a large network or from multiple independent networks. Conditional on a comprehensive set of controls, we assume independence across clusters, while allowing for complex, unrestricted dependence within each cluster. $\Box$
@@ 2.2 Exclusion restrictions
Assume that random regressors $x_{\ell}$ and errors $e_{\ell}$ have uniformly bounded second moments. There is a variety of exclusion restrictions one may impose in regression (1). The weakest of these is the contemporaneous exclusion restriction, $\mathbb{E}[x_{\ell}e_{\ell}]=0=\mathbb{E}[e_{\ell}]$ for all $\ell$, which defines the parameter $\beta$. When all observations are independent, this restriction (under mild regularity conditions) ensures that ordinary least squares (OLS) yields a consistent estimator of $\beta$. However, as we show later, in the presence of clustered data when the regression includes cluster fixed effects, imposing only contemporaneous exclusion fails to yield a consistent estimator of $\beta$.
At the other extreme is the strict exogeneity condition: $\mathbb{E}[e_{\ell}\!\mid\!x]=0$, which implies $\mathbb{E}[x_{\tilde{\ell}}e_{\ell}]=0$ for all $\tilde{\ell}$ and $\ell$. Under strict exogeneity, OLS is unbiased conditional on $x$. In particular, by the Frisch-Waugh-Lovell theorem, we have:
@@
where $y=[y_{1},\dots,y_{n}]^{\prime}$, $x=[x_{1},\dots,x_{n}]^{\prime}$, and $\tilde{x}=Mx$ denotes the residual from regressing $x$ on $W$. Conditionally on $x$, the regression errors $\{e_{\ell}\}$ are the only source of randomness. Since $\hat{\beta}^{\mathrm{LS}}$ is linear in the errors, standard results—such as unbiasedness, the law of large numbers, and the central limit theorem—hold. The only complication introduced by the clustered data structure is the need to adjust standard errors appropriately: cluster-robust standard errors are required because the relevant level of independence is at the cluster level.
Strict exogeneity is often too strong to hold in applications with clustered data, where unmodeled dependence between regressors and outcomes within clusters is likely. Researchers should tailor the exclusion restrictions to reflect institutional or structural features of the setting. We assume the researcher can specify and provide a matrix of exclusion restrictions, denoted by $\mathcal{E}$, an $n\times n$ indicator matrix encoding the imposed moment conditions. Specifically, $\mathcal{E}_{\tilde{\ell}\ell}=1$ encodes the assumption $\mathbb{E}[x_{\tilde{\ell}}e_{\ell}]=0$, while $\mathcal{E}_{\tilde{\ell}\ell}=0$ encodes the absence of a restriction and allows $\mathbb{E}[x_{\tilde{\ell}}e_{\ell}]\neq 0$.
@@
Suppose a researcher is interested in estimating the effect of program participation $x_{it}$ on student achievement $y_{it}$. It is plausible that current achievement $y_{it}$ could influence future eligibility or desire to participate $x_{is}$ for $s>t$, thereby inducing correlation between the current error term $e_{it}$ and future regressors $x_{is}$, violating strict exogeneity. In such settings, it is common to assume weak exogeneity: $\mathbb{E}[e_{it}x_{is}]=0$ for $s\leq t$, meaning the error is uncorrelated with current and past regressors but may influence future regressors. This structure arises naturally in dynamic panel models when, for instance, the lagged outcome is included as a regressor, $x_{it}=y_{i,t-1}$, or more generally when regressors and outcomes jointly evolve over time. In this case, the matrix $\mathcal{E}$ has triangular blocks of zeros if the observations are ordered by clusters and by time within each cluster.
If clustering is defined at the family level, the researcher may also be concerned about sibling feedback effects. For example, if the experience of an older child with the program affects a younger sibling’s participation. In this case, the researcher may refrain from imposing exclusion restrictions across such pairs, encoding this by zeros in the corresponding entries in $\mathcal{E}$, linking the older sibling’s error with the younger sibling’s regressor. $\Box$

### v2

@@ Abstract
We study linear regression models with clustered data, high-dimensional controls, and intricate exclusion restrictions. We propose a correctly centered internal instrument IV estimator that accommodates a broad class of exclusion restrictions and allows within-cluster dependence. The estimator admits a simple leave-out interpretation and is computationally tractable. We derive a central limit theorem for the associated quadratic form and propose a robust variance estimator. We also develop identification-robust inference procedures. Our framework extends dynamic panel methods to general clustered settings. We illustrate the approach in a large-scale fiscal intervention in rural Kenya, where spatial interference generates the exclusion-restriction pattern.
Keywords: weak exogeneity, many controls, interference, OLS inconsistency, internal instrument, leave-out approach
@@ 1 Introduction
Clustered datasets—such as panel, network, spatial, or group-structured data in which individuals or firms are observed repeatedly or nested within groups—have become increasingly common in empirical research. These settings present several methodological challenges. One is the need to accommodate heterogeneity, often addressed by including fixed effects and flexible time or group trends, which typically requires many control variables. A more fundamental difficulty is within-cluster dependence: observations in the same cluster may be correlated due to, e.g., spatial and network interference, spillover effects, and time-series dependence. Such dependencies complicate statistical inference and limit the effective aggregation of information across observations.
@@
The concept of exogeneity becomes more nuanced in the presence of clustered data. As we demonstrate, assuming only a per-observation exclusion restriction—uncorrelatedness of the error with the regressor within an observation—may deliver no identifying variation for consistent estimation of structural parameters. By contrast, strict exogeneity—each error term uncorrelated with all regressors in the cluster—is often implausible in many empirical contexts.
@@
To bridge strict exogeneity and unrestricted dependence, we assume the error term is uncorrelated with a subset of within-cluster regressors; the subset is application-specific. In panel data, it is common to assume errors are mean-zero conditional on current and past (but not future) regressors, allowing feedback from shocks to future policies. In spatial contexts, localized spillovers motivate analogous restrictions: regressors corresponding to sufficiently distant units within a cluster may plausibly be uncorrelated with a given error term, even if nearby regressors are not. Examples include spatial leakage in deforestation (jayachandran2017cash), spillovers in policing (blattman2021place), and fiscal intervention experiments in rural Kenya (egger2022general). In network settings, treatment effects may propagate along social ties, so regressors associated with nodes that are unconnected—or sufficiently distant in the network—may be uncorrelated with a given error term (paluck2016changing). These partial exogeneity assumptions more accurately reflect the dependence structures present in many empirical applications.
@@
When only a subset of exclusion restrictions holds, it becomes inappropriate to treat regressors as fixed or to condition on them. Their stochastic variation may correlate with the errors, undermining standard arguments that rely on conditioning on regressors. As a result, even estimators like OLS must be interpreted as ratios of stochastic quadratic forms, which introduces several challenges.
@@
The most well-known is Nickell bias (nickell1981biases), which occurs when the expected value of the OLS error numerator is nonzero, leading to asymptotic bias. While prominent in dynamic panels with fixed effects and lagged outcomes, analogous bias arises more broadly under clustered dependence. A second challenge is inference: standard variance estimators often fail to account for dependence among cross-cluster terms in the quadratic form. Finally, the denominator may remain asymptotically random, unlike in classical settings, leading to weak identification and further complicating inference.
@@
This paper studies the estimation of structural parameters in linear regression models with clustered data, high-dimensional controls, and researcher-specified exclusion restrictions. A key feature of the framework is the grouping of observations into disjoint clusters, permitting within-cluster dependence and independence across clusters. The setting accommodates unbalanced panels with multiple high-dimensional fixed effects, as well as spatial, network, and group-structured data. Our approach extends dynamic panel methods to a broader class of models, addressing the central challenges of bias, inference, and identification.
@@
Our first contribution is to characterize a class of correctly centered internal instrument estimators that remove the leading asymptotic bias of OLS. The approach accommodates high-dimensional exogenous controls and adapts to the exclusion structure specified in the application. The resulting estimator is an internal instrument IV estimator, defined as the one closest to OLS in a specified norm. The estimator is easy to implement and asymptotically efficient under some conditions.
@@
The procedure has a simple interpretation: for each observation, controls are partialled out using only those observations whose errors are uncorrelated with the observation’s regressor, yielding an observation-specific leave-out projection. A just-identified IV regression is then performed on the transformed equation, using the original regressor as the instrument.
@@
We further characterize the efficiency loss from varying the strength of exclusion restrictions. In particular, assuming exogeneity only at the observation level eliminates all identifying variation when fixed effects are present. This observation underscores the importance of carefully specifying the exogeneity structure in applied work.
@@
Our second contribution addresses inference. We show that, outside special cases such as (i) strict exogeneity and (ii) block-diagonal residualization (e.g., models with only cluster fixed effects), the estimator’s numerator is a nontrivial quadratic form in the errors. In these more general settings, standard cluster-robust variance estimators may fail, and valid inference requires central limit theorems for clustered quadratic forms. These issues are especially acute in models with many high-dimensional controls, such as two-way fixed effects (verdier2018estimation). We propose a new central limit theorem and variance estimator that account for this structure.
@@
Our third contribution addresses weak identification, which arises when the internal instrument—though valid—captures little identifying variation due to many controls or weak exclusion restrictions. As in the classical dynamic panel literature (blundell1998initial; bun2010weak), weak identification can compromise standard inference. We develop inference procedures that remain valid even when the estimator’s denominator exhibits substantial sampling variability.
@@
We propose a comprehensive empirical strategy that combines a correctly centered estimator with identification-robust inference and confidence sets. We apply it to a prominent randomized evaluation of a large-scale fiscal intervention in rural Kenya (egger2022general). A key challenge in this setting is spatial interference: treatment in one village affects outcomes in neighboring villages, complicating the choice of plausible exclusion restrictions. We show that weakening the exclusion restrictions yields less precise estimates and wider confidence sets.
@@
This paper contributes to several strands of the econometrics literature. It relates to the extensive work on linear dynamic panel data models—a mature field with too many important contributions to list exhaustively. For recent overviews on correcting Nickell bias, see the surveys by okui20211 and bun2015dynamic. Our estimation approach is more closely aligned with the internal instrument framework developed in anderson1981estimation, arellano1991some, ahn1995efficient, alvarez2003time, and blundell1998initial, and we extend the framework to accommodate more general models. In addition, we directly address weak identification concerns noted in the dynamic panel literature, particularly in bun2010weak. We also contribute to the emerging literature on quadratic central limit theorems for clustered data structures. These results have recently been employed in many-instrument panel settings, such as ligtenberg2023inference, and are crucial for understanding the distribution of estimators that are quadratic functionals of the errors.
@@
The remainder of the paper is organized as follows. Section 2 describes the data structure, a plausible set of exclusion restrictions, and two modeling perspectives—outcome-based and design-based. Section 3 characterizes correctly centered internal instrument estimators, introduces our proposed estimator as the solution to an efficiency optimization problem, and interprets it as a leave-out internal instrument procedure. Section 4 addresses uncertainty quantification, and Section 5 establishes a central limit theorem for quadratic forms with clustered data. Section 6 develops inference procedures robust to weak identification and a corresponding variance estimator. Section 7 illustrates the full strategy in an empirical application, showing how both the estimator and, especially, its uncertainty depend on the maintained set of assumptions.
@@
We use $0<c<1$ and $C>0$ to denote generic finite constants that may vary across equations but do not depend on the sample size. For a matrix $A$, we let $A^{+}$ denote its Moore–Penrose generalized inverse, $\operatorname{tr}(A)$ its trace, $\|A\|_{F}^{2}=\operatorname{tr}(A^{\prime}A)$ its Frobenius norm squared, and $\|A\|$ its operator (spectral) norm.
@@ 2.1 Data structure
where $\ell$ indexes the $n$ observations in the dataset. The parameter of interest, $\beta$, is a scalar, though most of our results extend to vectors of fixed dimension. The vector of strictly exogenous controls $w_{\ell}$ has dimension $K$, which may be large but is assumed to satisfy $K<n$. We denote by $W=[w_{1},\dots,w_{n}]^{\prime}$ the matrix of control variables. For simplicity, we treat $W$ as non-random; equivalently, all results may be interpreted conditional on $W$. Throughout, we assume that $W$ has full rank and define the projection matrix $M=I_{n}-W(W^{\prime}W)^{-1}W^{\prime},$ which projects out the variation associated with the controls $W$.
@@
Consider a regression with individual and time fixed effects: $y_{it}=\alpha_{i}+\mu_{t}+\beta x_{it}+e_{it},$ where $y_{it}$ and $x_{it}$ denote the outcome and treatment for student $i$ at time $t$. Let $\ell=\ell(i,t)$ index the observations, and define clusters as $S_{i}=\{\ell(i,t):t\text{ observed for individual }i\}$. Students may be observed across different, potentially non-overlapping periods. Here, following the standard panel-data setup, clusters coincide with individuals indexed by $i$. In other applications, however, clusters may be defined more broadly—for example, at the family level when students are siblings. ∎
@@
Suppose we study the causal effect of a municipality-level policy $x_{\ell}$ on outcome $y_{\ell}$. Assume that municipalities are nested within metropolitan areas indexed by $i$, which serve as the clusters. A rich set of controls is included—such as cluster fixed effects and federally influenced covariates—so that $\{(x_{\ell},e_{\ell}):\ell\in S_{i}\}$ is independent across clusters, but arbitrary dependence is allowed within clusters. ∎
@@
Consider treatment and outcome data for individuals connected via a social or informational network. The data may be drawn either from distinct clusters within a large network or from multiple independent networks. Conditional on a comprehensive set of controls, we assume independence across clusters, while allowing for complex, unrestricted dependence within each cluster. ∎
@@ 2.2 Exclusion restrictions
Assume that random regressors $x_{\ell}$ and errors $e_{\ell}$ have uniformly bounded second moments. There are a variety of exclusion assumptions one may make in regression (1). The weakest of these is the contemporaneous exclusion restriction, $\mathbb{E}[x_{\ell}e_{\ell}]=0=\mathbb{E}[e_{\ell}]$ for all $\ell$, which defines the parameter $\beta$. When all observations are independent, this restriction (under mild regularity conditions) ensures that ordinary least squares (OLS) yields a consistent estimator of $\beta$. However, as we show later, in the presence of clustered data, when the regression includes cluster fixed effects, the contemporaneous exclusion assumption alone fails to yield a consistent estimator of $\beta$.
At the other extreme is the strict exogeneity condition: $\mathbb{E}[e_{\ell}\!\mid\!x]=0$ where $x=[x_{1},\dots,x_{n}]^{\prime}$, which implies $\mathbb{E}[x_{\tilde{\ell}}e_{\ell}]=0$ for all $\tilde{\ell}$ and $\ell$. Under strict exogeneity, OLS is unbiased conditional on $x$. In particular, by the Frisch-Waugh-Lovell theorem, we have:
@@
where $y=[y_{1},\dots,y_{n}]^{\prime}$, $e=[e_{1},\dots,e_{n}]^{\prime}$ and $\tilde{x}=Mx$ denotes the residual from regressing $x$ on $W$. Conditionally on $x$, the regression errors $\{e_{\ell}\}$ are the only source of randomness. Since $\hat{\beta}^{\mathrm{LS}}$ is linear in the errors, standard results—such as unbiasedness, the law of large numbers, and the central limit theorem—hold. The only complication introduced by the clustered data structure is the need to adjust standard errors appropriately: cluster-robust standard errors are required because the relevant level of independence is at the cluster level.
Strict exogeneity is often too strong to hold in applications with clustered data, where unmodeled dependence between regressors and outcomes within clusters is likely. Researchers should tailor the exclusion restrictions to reflect institutional or structural features of the setting. We assume that the researcher can specify a matrix of exclusion restrictions, denoted by $\mathcal{E}$, which is an $n\times n$ indicator matrix encoding the moment conditions the researcher is willing to impose. Specifically, $\mathcal{E}_{\tilde{\ell}\ell}=1$ encodes the assumption $\mathbb{E}[x_{\tilde{\ell}}e_{\ell}]=0$, while $\mathcal{E}_{\tilde{\ell}\ell}=0$ encodes the absence of a restriction and allows $\mathbb{E}[x_{\tilde{\ell}}e_{\ell}]\neq 0$.
@@
Suppose a researcher is interested in estimating the effect of program participation $x_{it}$ on student achievement $y_{it}$. It is plausible that current achievement $y_{it}$ could influence future eligibility or desire to participate $x_{is}$ for $s>t$, thereby inducing correlation between the current error term $e_{it}$ and future regressors $x_{is}$, violating strict exogeneity. In such settings, it is common to assume weak exogeneity: $\mathbb{E}[x_{is}e_{it}]=0$ for $s\leq t$, meaning the error is uncorrelated with current and past regressors but may influence future regressors. This structure arises naturally in dynamic panel models when, for instance, the lagged outcome is included as a regressor, $x_{it}=y_{i,t-1}$, or more generally when regressors and outcomes jointly evolve over time. In this case, the matrix $\mathcal{E}$ has triangular blocks of zeros if the observations are ordered by clusters and by time within each cluster.
If clustering is defined at the family level, the researcher may also be concerned about sibling feedback effects. For example, if an older child’s experience with the program affects a younger sibling’s participation. In this case, the researcher may refrain from assuming exclusion restrictions across such pairs, encoding this by zeros in the corresponding entries in $\mathcal{E}$, linking the older sibling’s error term with the younger sibling’s regressor. ∎
Example 2 (continued).
Suppose a researcher is interested in estimating the average treatment effect of a policy using a randomized controlled trial, in which an economic resource is randomly assigned across municipalities $\ell$. Spatial interference arises when outcomes in a given location may also be affected by treatment in neighboring locations—for example, because of resource sharing or price adjustments in shared local markets. Such spillover effects violate the exogeneity assumption across neighboring localities, since treatment in one location directly affects outcomes in the neighboring locality. If the researcher believes that interference occurs only when locations are within $R$ kilometers of each other and is negligible at larger distances, then the matrix $\mathcal{E}$ encodes indicators of whether two locations lie more than $R$ kilometers apart. Notable examples of such settings are spatial leakage in deforestation (jayachandran2017cash), spillovers in policing (blattman2021place), or a fiscal intervention experiment in rural Kenya (egger2022general). ∎
Example 3 (continued).
Consider the randomized controlled trial studied in (paluck2016changing), in which students are randomly assigned to participate in a program that teaches teenagers conflict-resolution skills. A researcher may believe that the treatment affects not only the outcomes of treated students, but also those of their close friends. In this case, it is reasonable to assume that an exclusion restriction holds only for pairs of observations that are not directly connected by friendship, so that the matrix $\mathcal{E}$ is given by the complement of the friendship adjacency matrix. Such spillovers of information or resources through social networks are increasingly common in modern randomized controlled trials. Prominent examples include banerjee2013diffusion, banerjee2019using, and banerjee2024can. ∎